OpenStudy (anonymous):

1 2 ---- - ------ = 2 3x+9 x+3

OpenStudy (anonymous):

3x+9<----you can common factor the 3 out so you get3(x+3)

OpenStudy (anonymous):

since your subtracting the denominator as to be the same

OpenStudy (anonymous):

answer choices. A. x=1 B. x=-23/6 C. x=-5/2 D. x=-5

OpenStudy (anonymous):

okay ..

OpenStudy (anonymous):

i agree with @nincompoop i got x=-23/6 as well

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

\[\frac{ x+3-6x-18 }{ 3x^2+9x+9x+27 }=2=\frac{ -5x-15 }{ 3x^2+18x+27 }\] 6x^2+36x+54=-5x-15 6x^2+41x+69=0 (x+3)(6x+23)=0 x=-3 or x= -23/6